Weather: Initially calm and warm, then windy and overcast in the afternoon

Staff: Just me and Gary.  Justin is taking a well-deserved day off.

There were only 2 items on our punch list: fixing the t7 sonic and the 28m ec150/ec100 config.

I spent most of the day working on qctables and webplots and now have all running as I intended – not to say we can't add more plots.

The students arrived back from yesterday's trip to Vegas with a 1000W pure sine inverter, which Chenning tested and worked fine.  They installed it (though deciding to connect directly to the battery, rather than the Victron load output) and it seems to be running.  I may still add a solar panel or two to his setup.  (Just trying to decide if one should be added to t0, before committing more to Chenning.). I'm still not sure if he has the DTS fiber running or not.

The students (with Bill) launched 2 sondes today.

Sounds like everyone was somewhat happy with the way things are working for our first day.  I certainly am.

Tomorrow's tasks: